Solve the system by addition.
x + y = 6
x – y = 4

Answer by rapaljer(4671)

Just add the two equations together:
x + y = 6
x – y = 4
2x=10
x=5

Now, substitute this value of x back into either equation (the first one will do!)
x+y=6
5+y=6
y=1

Check by substituting both numbers into the second equation:
x-y=4
5-1=4 It checks!!

R^2 at SCC